OpenStudy (anonymous):

A ball is released from a hot air balloon moving downward with a velocity of -10.0 meters/second and a height of 1,000 meters. How long did it take the ball to reach the surface of Earth? Given: g = -9.8 meters/second2.

OpenStudy (anonymous):

what formula do i use for this equation

OpenStudy (anonymous):

h - ho =1/2g(t-to)^2 + vo(t-to)

OpenStudy (anonymous):

what

OpenStudy (anonymous):

find t? from equation above

OpenStudy (anonymous):

how

OpenStudy (anonymous):

when to = 0,ho=1000m, vo=-10m/s

OpenStudy (jt950):

S=ut+(1/2)gt^2

OpenStudy (jt950):

1000=(-10)(9.8)+(1/2)(9.8)(t^2) t=14.97s

OpenStudy (anonymous):

ok so how would i work that

OpenStudy (anonymous):

look direction

OpenStudy (anonymous):

-1000= - + (-)

OpenStudy (anonymous):

confusing

OpenStudy (anonymous):

g and v the same direction

OpenStudy (anonymous):

-1000=-1/2gt^2 _ 10t